40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 4 of 57  Not logged in ELOG logo
ID Date Author Type Category Subject
  2719   Wed Feb 9 15:21:27 2022 ranaHowtoTempCtrlplots

In the plots comparing data and models, can you use the legend to indicate which is which? e.g. use dots for data and solid lines for models, and then label them as that in the legend. Also nice to include error bars on the temperautre measurements; I think there's a python way to plot this as a shaded region as well.

  2718   Mon Feb 7 16:06:36 2022 StephenDailyProgress 31 Jan Fastest Radiative Cooling run, ended 07 Jan

[Radhika, Stephen]

The heater was turned on at 3:13pm on Friday 2/4.

We specified a set temperature of 123K. However, the CTC100 PI control included a 1 W lower limit on the input to the heater, so there was a steady load of 1 W applied to the Silicon Workpiece over the weekend.

At 16:01 the cryocooler was turned off to start the warmup.

The CTC100 PI control was configured with a setpoint of 250 K on the Workpiece RTD, to aid in the warmup, and an allowable power range from 0 W to 22 W.

  2717   Mon Feb 7 11:57:38 2022 StephenDailyProgressCryo vacuum chamberDoubled Thermal Linkage Capacity run started 11 Feb 2022

[Stephen, Radhika]

The double copper bar configuration pictured in Attachment 1 has been implemented. We completed the updates within work sessions on Thursday and Friday. Here's a pseudo-log:

  • Removed all items from coldplate to make space for thermal linkage.
  • Prepared thermal linkage for installation.
    • ref. QIL/2694 for all preparations made.
  • Installed thermal linkage.
    • Removed thermal linkage to allow for regreasing
      • Discovered a single brass #4 washer sandwiched between coldhead and rigid copper bar (likely lost during efforts to install an RTD during intitial thermal linkage install); thermal grease was mostly indented at brass washer and at area opposite location of brass washer, suggesting non-uniform contact.
    • Applied Apiezon N thermal grease at joints: [coldhead - top copper bar], [top copper bar - bottom copper bar, near coldhead], [top copper bar - bottom copper bar, near thermal strap].
    • Sandwiched together top and bottom copper bars, to insert simultaneously.
    • Angled top and bottom copper bars to pass below coldhead, and overall passing the two bars in together went smoothly.
    • Bolted both copper bars into coldhead using long (1.25") 4-40 bolts threaded into coldhead. Access was from bottom port below cryocooler. Torqued using short arm of allen key (no access for long arm).
    • Borrowed one bolt location for RTD spring clamp mount. Confirmed continuity of coldhead RTD.
  • Repaired pins damaged in removal, and inserted sheaths intended to replace the previous scheme and hopefully prevent pin and socket damage. More detail later :).
  • Installed shields.
    • Applied new Apiezon N thermal grease to bottom flange of inner shield.
    • Re-clamped inner shield to cold plate using a dog clamp array.
    • Placed outer shield in contact with outer shield; rocking percieved when pressing on one quadrant of the annulus.
    • Confirmed continuity of shield RTDs.
  • Installed test mass, still with heater mounted to planar surface as in QIL/2715.
    • Confirmed suspension wires were contacting bare silicon OD surface in middle.
    • Confirmed continuity of heater and of workpiece RTD.
  • Closed up, pumped down, cooled down.
    • Roughing pump on at ~4:20 pm
    • Turbo on with no issues after 15 minute programmed delay
    • Cooling started at 5:15 pm, pressure was a few E-4 torr; no need to wait so long, that was just when we got around to it.

All images are (or soon to be) posted to the QIL Photo Dump.

Attachment 1: D2000310_y-004_section_double_linkage_20220207.png
D2000310_y-004_section_double_linkage_20220207.png
  2716   Fri Feb 4 14:00:19 2022 RadhikaDailyProgress 31 Jan Fastest Radiative Cooling run started

Attached are best fits for cooldown runs on 01/14 and 01/31. The setup for both cooldowns can be found in the previous ELOGs. We noticed that the outer shield did not cool as significantly on 01/31 than on 01/14, hinting that there might have been more thermal contact between the outer shield and cold plate / copper bar. 

The model considers the resistances of the following conductive elements (and uses these resistances as fit parameters):

   1.) Joint between cold head and copper bar
   2.) Bulk of copper bar
   3.) Joint between copper bar and flexible strap
   4.) Bulk of flexible strap
   5.) Joint between flexible strap and cold plate

These additions helped the model more closely resemble our recorded data, with a few exceptions:

   - At early cooldown times, the model seems to be underestimating the heat load on the inner shield and outer shield.

   - The best fit was performed on the inner shield and outer shield data (to fine tune elements of the cold linkage), so the test mass fit is not optimized. (This will be performed next to refine emissivity predictions.)

To identify bottlenecks in the cold linkage, I used the 01/31 model and tweaked the resistances to see which would provide the largest gains in cooldown. The results from such tweaks are below:

   - Halfing the resistance of the greased joints (1, 3, 5) made negligible change to the cooldown. 
   - Halfing the resistance of the bulk of the copper bar made significant improvement to cooldown (Attachment 3).
   - Halfing the resistance of the bulk of the flexible strap made some improvement to cooldown (Attachment 4).

From these observations, it seems like the greased joints are thermally efficient, and the bulk area of the copper bar appears to be the largest bottleneck. 

 

Attachment 1: Cooldown_0114_analyzed.pdf
Cooldown_0114_analyzed.pdf
Attachment 2: Cooldown_0131_analyzed.pdf
Cooldown_0131_analyzed.pdf
Attachment 3: Cooldown_0131_copperbar_halfR.pdf
Cooldown_0131_copperbar_halfR.pdf
Attachment 4: Cooldown_0131_flexstrap_halfR.pdf
Cooldown_0131_flexstrap_halfR.pdf
  2715   Mon Jan 31 15:40:21 2022 StephenDailyProgress 31 Jan Fastest Radiative Cooling run started

As discussed during the 21 Jan 2022 meeting, the next cryostat run will seek the fastest radiative cooling (again, see QIL/2706) through the following configuration choices:

  • move heater to test mass
  • add thermal grease to inner shield - coldplate interface (clamped joint)

Actions completed 27, 28, and 31 Jan 2022

  • Vent.
  • Remove Test Mass in frame.
    • For simplicity, we snipped the RTD lead (we didn't want to have to redo the cryovarnish joint).
      • Of course, the joint failed during the repairs so we had to redo the workpiece RTD cryovarnish joint anyway!
  • Use cryovarnish to affix heater.
    • Removed aquadag in area of joint using IPA-soaked Alpha wipe and scrubbing motion.
    • First, varnished cigarette paper anchor pad, then added heater.
    • We allowed the cryovarnish to cure overnight between the 27th and 28th, then allowed the joint to hang in its vertical orientation over the weekend to confirm its integrity.
  • Use conventional pin-socket arrangement to add junction to heater leads and workpiece RTD leads.
    • Confirm all RTD leads are functional, and repair crimp joints and Kapton tape insulation where needed.
  • Clean up aquadag flakes on coldplate.
  • Remove inner shield, add Apiezon M thermal grease, and bolt down again.
  • Touch up position of mylar shield on cold linkage, aluminum foil aperture covers.
    • All apertures are covered except:
      • Inner Shield: cold linkage (not an issue - partially occluded by copper bar, mylar tube extends into cold head view and mylar mitten)
      • Outer Shield: electrical port, cold linkage (aperture has been extended to have additional area underneath the original circular aperture)
  • Insert Test Mass in frame. Dog clamp down frame.
  • Pump down, cool down.
    • Pump down started at ~15:15.
    • Cool down started at 15:44.

Model updates required to reflect new configuration:

  • estimate conductive leak from 24 AWG heater leads to test mass
  • model conductance of cryovarnish joint between heater and test mass - does cigarette paper make a difference?
  • model greased joint of inner shield - cold plate interface
    • confirm clamp load is adequate for effective greased joint, at room temp and cold temp!
  2714   Fri Jan 28 10:31:15 2022 ranaSummaryTutorial videooh no, stap the units madness, aaaaahhhh!!! noooooo!!!!

What if I told you SI-units is way harder to miscalculate with - Matrix  Morpheus | Make a Meme

 

  2713   Tue Jan 25 14:01:27 2022 StephenDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

There's also a convention to write "50 kgf" to designate "kilograms of force" (implying the same conversion Rana describes, multiplying by g). I see kgf enough in the mechanical engineering world that I wouldn't have been confounded, so I wanted to pass that along.

Quote:

I think its least confusing to just replace 50 kg g with 500 N. Writing 50 g can be misleading, it seems like 50 grams.

 

 

  2712   Tue Jan 25 13:15:40 2022 ranaDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

I think its least confusing to just replace 50 kg g with 500 N. Writing 50 g can be misleading, it seems like 50 grams.

 

  2711   Tue Jan 25 10:24:14 2022 ranaDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

For one surface resting on another, the force of the contact is the grativational force of the top surface. There's an implicit factor of g that cancels out from the ratio, so it becomes a ratio of masses. The heat conductance listed for a 50 kg Al-Al pressure joint is interpreted as the heat conductance for a force of 50g.

Quote:

Is there a units issue? 50 kg is a mass, but not a force.

 

  2710   Fri Jan 21 17:32:40 2022 ranaDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

Is there a units issue? 50 kg is a mass, but not a force.

  2709   Fri Jan 21 13:28:48 2022 RadhikaDailyProgressCryo vacuum chamberHeat transfer between grease joints and pressure joints

I took a deep dive into Ekin 2.6 to understand heat tranfer between various joint types, specifically pressure joints and grease joints. This was motivated by the fact that modeling of the cold plate and inner shield temperatures seemed to be missing some key physics. 

Heat conductance through grease joints is dependent on the contact area of the surfaces:

\frac{\dot{q}}{\Delta T} = \frac{\dot{q}}{\Delta T}(1 cm^2) \frac{A}{1 cm^2}, where \frac{\dot{q}}{\Delta T} = the heat conductance. The heat conductance of a grease contact with area 1 cm2 can be found in Attachment 1, indicated by the region between the blue and red boxes (our temperature range).

Heat conductance through pressure joints is dependent on the force holding the surfaces in contact:

\frac{\dot{q}}{\Delta T} = \frac{\dot{q}}{\Delta T}(50kgf) \frac{F}{50kgf}, where the heat conductance of a pressure contact with force of 50kgf is found in Attachment 1, indicated by the orange box. (Megastat pressure contacts are between the cold plate and shields, so the Al-Al contact was referenced.)

[Added] Heat conductance through varnish joints are similar to grease joints, in that the conductance scales with contact area of the joint. (It follows the same formula as above for grease joints.) The heat conductance of a varnished contact with area 1 cm2 can be found in Attachment 1, and note that it is higher than that of greased contacts.

My efforts this week went into incorporating these heat links into the model. This required splitting up components into various parts (verging on a finite-element approach), since every joint is considered in addition to the elements themselves. A few notes:

1. For now, I assume the force between the inner shield and cold plate is 50 kg, for simplicity. Therefore, the heat conductance being used for this pressure joint is 4.5 W/K (from chart), and I am using this constant value across temperatures until a better estimation can be found.

2. For the grease joints, I estimate the heat conductance at 1 cm2 to be ~ 1 W/K, for the 50-300K range.

3. The contact between the outer shield and cold plate is not actually uniformly touching, as noted in 2706. I am not sure how to estimate the actual force between the surfaces, so I will add this as a fit parameter in the model. 

The new model with this incorporation still needs some adequate debugging, but I felt these were vital steps to ensure we get realistic answers regarding cooling power of the copper bar vs. LN2. Once I feel the model can be trusted, I will follow up with analysis of the new optimized cooldown. 

Attachment 1: thermal_conductance_joints.png
thermal_conductance_joints.png
  2708   Tue Jan 18 16:11:35 2022 StephenDailyProgressCryo vacuum chamberStatus of Fastest Radiative Cooling run started 14 Jan

I ended the first cooldown of the run at 12:49 today, approximately 92 hours after it had started. Workpiece temperature was 100 K - more data and model fitting will reveal more insights on the results.

I started a 5 W injection through the heater (dog-clamped to baseplate with indium gasket), which I intend to leave running until steady state temperature is reached, at least overnight. This power level will not present a risk to the system, so I feel comfortable with this static power input even though it is not interlocked.

Update after 25 hours [ATTACHMENT 1]

The 5 W heating of the shields has not quite leveled off yet, and the workpiece temperature resumed falling, showing the temperature had not yet reached a steady state. I will leave 5 W on until I see a steady state, then I will plan to turn off 5 W and allow the workpiece trend to level off fully. Summary:

  • with no heater input, after 90+ hours the workpiece was slowly cooling at a rate of ~ 1 K per 6 hours, and the inner shield temperature was stable at about 76 K.
  • with 5W heater input, after about 25 hours the workpiece is (still) slowly cooling at a rate of ~ 1 K per 18 hours, and the inner shield temperature is about 88 K and slowly rising about 0.1 K per 7 hours

See attachment 1 for data viewer screenshot which reflects the above summary.

Update after 145 hours [ATTACHMENT 2]

Continued injecting 5 W for the whole weekend. Over the last 24 hours, the workpiece temperature seems to have leveled off into a new cooling rate, while both shields alternated between heating and cooling. I think this can be described as a steady-enough state that I'm ending the cooldown.

Now, I'll let the chamber come up to temperature with the help of the 5 W load.

Attachment 1: 5w_input_qil_cryostat_20220119.png
5w_input_qil_cryostat_20220119.png
Attachment 2: 5w_input_121_to_145_hours_qil_cryostat_20220124.png
5w_input_121_to_145_hours_qil_cryostat_20220124.png
  2707   Mon Jan 10 16:46:19 2022 PacoMiscEquipment transfer2 inch GYRO optics

[Paco, Yehonathan]

This afternoon, we entered the QIL to look for a 2 in optic from the old GYRO experiment that might work as the PR2 in the BHD upgrade in 40m. We found some in the right most cabinet near where MEGASTAT is. We ended up taking two pieces (see Attachment #1 for evidence) one labeled "COATED 2"dia F.S. SUB S1: HR R>99.99% @ 1064NM S2: AR R<0.15% S/N:" and the other one from ATFilms, labeled "2"dia x 0.375" thk FS HR/AR @ 1064nm Run#s': V6-704 & V6-705" and a datasheet which specifies T = 0.02 % or 200 ppm.

Attachment 1: PXL_20220111_004553530.jpg
PXL_20220111_004553530.jpg
  2706   Fri Jan 7 14:54:58 2022 StephenDailyProgressCryo vacuum chamberFastest Radiative Cooling run started 14 Jan

As discussed during the 07 Jan 2022 meeting, the next cryostat run will seek the fastest radiative cooling through the following configuration choices:

  • eliminate radiation leak apertures
  • maximize emissivity of test mass volume
  • improve conductivity to outer shield

Actions completed 14 Jan 2022

  • Vent
  • Cover up aperture to viewport with foil (on both shields).
  • Remove foil from inner shield inner surface.
  • Place outer shield directly onto cold plate (no clamping force).
    • A bit unsatisfying, because the outer shield is not flat relative to the cold plate, but the shield is resting quite firmly with only a slight wobble when pressed. As with any plane-plane contact, there are a limited number of true points of contact, and the contact area is increased somewhat by foil that is folded around the outer shield or the cold plate.
      • I would estimate the contact area at ~5% of the total area of the bottom flange.
    • 2x G10 spacer washers were dropped into the lower volume whild making this switch.
    • Note that bottom lid of outer shield is still not conductively cooled, and therefore will be approximately room temperature.
  • Tidy up aluminum foil collar around bottom lid of outer shield.
  • Tidy up mylar shield around conductive link after inner shield
  • Pump down, cool down.
    • Pump down started at ~15:35.
    • Cool down started at 16:10.
  2704   Tue Dec 21 15:33:39 2021 StephenDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with worse inner shield inner surface emissivity - next run repeating

I opened the QIL cryostat today for a health check and visual inspection before the next run. Because I saw a couple of interesting issues, I decided to redo the same run with more attention to detail on the closeout. I'm worried the outer shield may have been linked to the inner shield and baseplate enough to affect comparability with the prior run.

issues with this  run, requiring redo:

  • RTD wire for outer shield was clamped under lid of inner shield - this might have created a conductive link between the inner shield and the outer shield
  • outer shield was more wobbly than usual, and could have possibly been off of the three spacers - this might have created a conductive link between the outer shield and the baseplate

run details:

  • pumping started at 3:45 pm
  • turbo started with 15 minute delay at 20 torr
  • cryocooler started at 4:15 pm with active ion gauge pressure at 3 e-4 torr.

And since I didn't get to implement any of the intended next runs, here are some notes on other future runs of interest:

  • Si mass covered with Al foil (matte side facing out) - interested in making the emissivity of the test mass equal to that of the inner shield in the new config, for modeling.
    • (of course, this emissivity equivalence would be an approximation, as there is a large area of the test mass which is bare silicon)
  • outer shield clamped/resting on baseplate - this is predicted by Koji to be the most efficient cooling configuration.
  • shielding attached to structure holding Si mass (electropolished aluminum, aquadag aluminum, bare aluminum surfaces are all available.
  2703   Thu Dec 16 17:57:15 2021 Radhika, StephenSummaryCryo vacuum chamberMegastat geometric parameters

This ELOG serves as a compilation of known/measured geometric parameters of Megastat. This is informative for thermal modeling of the system, so I wanted to create a centralized reference. A reference to these dimensions has been added to the Wiki page

Chamber specs
   Outer Radius = 0.3048 m (12")
   Wall thickness = .00477 m (.188")
   Height = 0.3048 m (12")
   Flange thickness = .0254 m (1")

Outer shield specs
   Outer Radius = 0.2794 m (11")--> CAD .265 m (10.433")
   Height = 0.2286 m (9") --> CAD .206 m (8.110")
   Thickness = 2.90 mm (0.114") (CAD nominally 3 mm, but 9 gauge aluminum is standard)

Inner shield specs:
   Outer Radius = 0.244983 m (9.645") --> CAD .225 m (8.858")
   Height = 0.205994 m (8.11") --> CAD .192 m (7.559")
   Thickness = 2.90 mm (0.114")

Cold plate specs:
   Radius = CAD .254 m (10")
   Thickness = CAD .01498 m (.5897")

Test mass specs: (confirmed)
   Radius = 0.0508 m (2")
   Length = 0.1016 m (4")

Copper bar specs:
   Length = 0.508 m (20") --> note that center to center length is .440 m in CAD
   Width1 = 0.03175 m (1.25") (bulk cross section, could be approximated accross full length)
   Width2 = 0.049784 m (1.96") (cross section at cold head bolting interface)
   Thickness = 0.011684 m (0.46")

Coldhead specs:
   Radius = 0.03175 m (1.25")
   Thickness = .0516 m

CAD (.EASM) is located at https://caltech.app.box.com/folder/131056505764 (File path: Voyager > Mariner > CryoEngineering). Screenshot of current state is added as Attachment 1.

CAD (source file, .sldasm - SolidWorks 2021) may be accessed via the PDM Vault (File path: llpdmpro > voyager > rnd qil cryostat)

Attachment 1: D2000310_y-003_20211222.png
D2000310_y-003_20211222.png
  2702   Thu Dec 16 15:54:44 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with worse inner shield inner surface emissivity - CTC100 temperature control success

This post will host plots and trends from this radiative cooling run. At a glance, the tuned CTC100 PI control was able to control the workpiece steady state temperature in this radiative cooling test within .005 K.

Run description: At 4 pm Wednesday, the workpiece temperature was at steady state from the QIL/2701 cooldown, a little less than 120 K. From 4pm Wednesday thru 5pm Thursday (25 hours) the CTC100 controller was actuating on the workpiece RTD temperature (cryovarnished to the suspended Si mass) using the resistive heater (dog clamped to baseplate with indium foil gasket). The conductive heating of the cold plate, and therefore the inner shield, led to radiative heating capacity (via ΔT)  that actuated on the temperature of the suspended test mass. As found in QIL/2643, the suspended Si mass is well isolated from conduction to the cold plate.

Before the run, the CTC100 PID controller was allowed to autotune using a long lag (600 s) and a moderate acutation step (10 W). After autotuning, the D term was still 0, which seemed fine.

Data: Attachment 1 plots cooldown curves for all RTDs during this run. Attachment 2 compares this run's test mass and inner shield temperature curves to those from the previous run (Aquadag on inner surface of inner shield). The expected result of this change (coating inner surface of inner shield with Al foil) is a weakened radiative coupling between the inner shield and test mass, leading to less effective cooling of the test mass. 

Initial observations from data:

1) The cold head temperature curve again suggests 2 time constants, and cooldown is identical.

2) The inner shield's cooldown is roughly unchanged.

3) The outer shield's temperature drops significantly more, indicating a stronger coupling to the inner shield. We will check for a conductive short the next time we open up.

4) The test mass's cooldown matches expectations (weaker radiative coupling).

[WIP - The data will be fitted and discussed]. More detailed analysis from fit to come, including from heater runs.

 

Attachment 1: cooldown_12-10_all.pdf
cooldown_12-10_all.pdf
Attachment 2: cooldown_12-10_vs_11-16.pdf
cooldown_12-10_vs_11-16.pdf
  2701   Fri Dec 10 15:58:57 2021 StephenDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with worse inner shield inner surface emissivity

Started a new run this afternoon, with the following goals:

   1) confirm that the first run (QIL/2695) went smoothly, by performing a visual inspection in the chamber while setting up for the first run.

      - kapton tape affixing inner shield RTD lead junctions to inner shield had fallen. These junctions were simply hanging - not ideal, but apparently not too harmful. Not likely to impact temperatures, in my opinion, but could have led to shorts or glitches in data.

      - all RTDs appeared to be fixed and well-contacted to surfaces

      - Everything seemed to be in good shape with the copper bar, no apparent issues

   2) obtain a second run with similar configuration, now that the rigid copper bar linkage has been implemented.

   3) vary a single important parameter relative to the first run, namely the inner surface emissivity of the inner shield, so that the impact of that parameter may be observed.

      - Added Aluminum foil (matte side visible) to the inner shield inner surfaces (lid and cylinder, both). Anywhere there was previously black Aquadag, there is now matte aluminum foil.

      - Kept the same apertures for viewport access and for electrical and thermal connection passthrough, basically attempted to achieve identical shield coverage.

      - There is one small sliver of black aquadag visible at the location of the electrical leads, but I didn't worry about patching that small area.

Run Details:

   - Pumps on at  ~3:40 pm

   - Cooling started at 4:13 pm (pressure ~6 mTorr, rapidly falling with turbo pump spinning up from ~70% to ~85% over a 1 minute interval). Coldhead RTD is responsive.

   - All photos will be posted to the QIL Cryo Vacuum Chamber photo album.

   - Note from check in on Monday afternoon, ~ 69 hours after start: everything looks good, and the workpiece temperature (~127 K) seems to reflect the emissivity change.

  2699   Sun Dec 5 17:33:50 2021 ranaDailyProgressCryo vacuum chamberTransitioning MS cooldown fitting to MCMC scheme

thats looking good

You should try to use either the corner or getdist packages to plot the 'corner' plots commonly used when showing correlated posteriors (cf. https://emcee.readthedocs.io/en/v2.2.1/user/line/) so that we can see what's up with the other uncertainties

Quote:

This week I have been working towards a Markov-chain Monte Carlo (MCMC) approach for fitting Megastat cooldown data and obtaining estimates on various emissivity values. I started with a simple model, only simulating the radiative cooling from the inner shield to the test mass. I supply the test mass and inner shield temperature data, and the emissivity of Aquadag (coating both the TM and inner surface of IS) as the only fit parameter. I am using Stan for the modeling, and Attachment 1 is a copy of my stan model.

The results of the simple model are in Attachment 2. The emissivity of Aquadag is estimated as a gaussian centered around 0.7. I am still determining whether this result is "real", or if the simulation is simply returning my prior. I will look into this more before adding complexity to the model.

 

  2698   Fri Dec 3 09:31:40 2021 RadhikaDailyProgressCryo vacuum chamberTransitioning MS cooldown fitting to MCMC scheme

This week I have been working towards a Markov-chain Monte Carlo (MCMC) approach for fitting Megastat cooldown data and obtaining estimates on various emissivity values. I started with a simple model, only simulating the radiative cooling from the inner shield to the test mass. I supply the test mass and inner shield temperature data, and the emissivity of Aquadag (coating both the TM and inner surface of IS) as the only fit parameter. I am using Stan for the modeling, and Attachment 1 is a copy of my stan model.

The results of the simple model are in Attachment 2. The emissivity of Aquadag is estimated as a gaussian centered around 0.7. I am still determining whether this result is "real", or if the simulation is simply returning my prior. I will look into this more before adding complexity to the model.

Attachment 1: MS.stan
functions {

  real coupling(real e_inner, real e_outer, real A_inner, real A_outer) {
    
    real E;
    E = e_inner*e_outer / (e_outer + (A_inner/A_outer)*(e_inner - e_inner*e_outer));
    return E;
  }
  
  real Cp_Si(real T) {
... 66 more lines ...
Attachment 2: stan_blackcoat_fit.pdf
stan_blackcoat_fit.pdf
  2697   Fri Nov 19 14:01:40 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with better shield emissivity

[WIP]

11/16 was the first Megastat cooldown after exchanging the copper braid linkage for a copper bar. Attachment 1 compares the cooldown trends for the test mass, inner and outer shields, and cold head. The solid curves are the new cooldown trends (copper bar), and the faded dashed curves are the previous cooldown trends (copper braid).

Immediate observations:

    - The coldhead has a reduced heat load, and interestingly a second time constant governs cooldown from ~2-35 hours.

    - The inner shield time constant is reduced significantly, but the inner shield experiences a slightly greater heat load at steady state.

    - The test mass cooling is improved as expected, given inner shield cooldown.

    - The coupling between the outer shield and inner shield has increased, resulting in greater cooling of the outer shield. This could explain the added heat load to the inner shield. 

Attachment 1: comp_cooldown_728_cooldown_1116.pdf
comp_cooldown_728_cooldown_1116.pdf
  2696   Fri Nov 19 09:36:52 2021 AnchalMiscEquipment transferBorrowing two 2" optics

I came and looked around for a Y1S HR coated 2" optic. I found two ATFilms labeled optics with information of only their substrate. The coating info is encoded in the run# but I could not find a place for what it means. So I'm gonna take these and measure the transmission in 40m.

  • 2"x.375" thk FS, HR/AR @ 1064, Run #'s: V6-704 & V6-705
  • 2" x 0.375" PL/PL Run #'s: V2-2239 & V2-2242
Attachment 1: PXL_20211119_174146457.jpg
PXL_20211119_174146457.jpg
  2695   Fri Nov 12 14:31:38 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with better shield emissivity

In this phase, we are working toward improving our setup with a rigid copper bar, and obtaining a new data point for our radiative cooling thermal models for a suspended silicon mass. Since the past cooling runs of a silicon test mass did not yet incorporate aquadag-painted shields, we wanted to obtain a new data point in the model (in other words, we painted the shields in QIL/2645, but the next test was a PD measurement, so this is the first silicon test mass measurment after shields were painted). The improvement to the thermal linkage, now using a rigid copper bar with higher conductivity (ref. QIL/2666), is a second variable being changed simultaneously in the spirit of improving the cooldown time.

Refer to the prior post (QIL/2694) for the bulk of the blow-by-blow of configuring the chamber to use the rigid copper bar linkage. This post will describe the mounting of the Si mass, and the pump down and cool down.

  • The silicon mass with Aquadag barrel was dropped into the existing frame, with the previous wire arrangement and with no particular requirement on position or orientation (just best effort centering and leveling). Adjustments were done chamberside as access was easier.
  • The frame was lifted into the chamber, with the hanging mass supported by auxiliary fingers, and placed in an available area. Since conductive cooling was not a dominant mode of heat transfer in this setup (ref. QIL/2647), clamping to the baseplate was simply a single dog clamp on each foot of the frame.
  • The cigarette paper was cryovarnished to the surface in the bare central position. Once the cryovarnish was set, the RTD was cryovarnished to the cigarette paper pad. No  strain relief or thermal anchoring considerations were implemented. RTD continuity was verified.
  • Lids were bolted down and shields were finalized (avoiding shorting to copper bar, making sure foil drapes covering apertures were well positioned, etc.
  • Vacuum pumps on at ~3 pm, cryocooler on at 3:30 pm. At 4 pm, things are still looking good!

Closeout photos will be posted to the QIL Cryo Vacuum Chamber photo album.

Attachment 1: IMG_2738.jpeg
IMG_2738.jpeg
Attachment 2: IMG_2741.jpeg
IMG_2741.jpeg
Attachment 3: IMG_2742.jpeg
IMG_2742.jpeg
Attachment 4: IMG_2745.jpeg
IMG_2745.jpeg
Attachment 5: IMG_2747.jpeg
IMG_2747.jpeg
  2694   Fri Nov 12 14:21:32 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberUpgrade to Rigid Copper Bar, and assorted transitions from PD testing

This post describes upgrade efforts from 10 - 16 November, with the following goals:

 - introducing a solid copper bar thermal linkage

 - shifting the setup away from PD testing

 - preparing for the next test (radiative cooling of Si)

Here are some highlights of the effort:

  • While removing the shields we found contaminants had plated near the line of sight surfaces at the optical window and the electrical feedthrough [Attachment 1]. The film was removed by IPA wipe [Attachment 2] and was not evident in any other location (presumably these were the cold surfaces in line of sight, so they received the most contaminant, but there may be a thinner deposition throughout!).
  • In order to install the copper rod, we needed to cut out slots in the outer shield and inner shield. We used a reciprocating saw and held the piece stiffly on a table [Attachment 3] [Attachment 4].
    • We tried to use large snips, but that failed to provide enough cutting force, especially where it was necessary to use the tip to access into the flanged corner. We also damaged a few of the electrical leads to the feedthrough (formerly used to wire the OSEMs) - this will require attention at next opportunity.
  • The copper rod itself was found to have a few issues:
    • Outer surfaces were somewhat tarnished and greasy, so an 80 grit aluminum oxide paper was used to clean up all surfaces [Attachment 5]. The surfaces were then wiped with IPA using Alpha wipes.
    • Slot width was matching the long thermal strap instead of the short thermal strap, so a drill press was used to add cut away the correct areas of one pair of holes [Attachment 6]. Later, this modification required larger washers in the stack under the bolt head.
    • The cold head had a larger OD than designed, so we were unable to use the corner holes to mount the RTD. There was not enough space at the corners to host the nut or washer. Instead one of the bolt pattern holes was used to host the RTD stack [Attachment 7]
  • The installation of the copper rod required the following steps:
    • We documented the previous state of the table [Attachment 8].
    • We removed the inner shield, the outer shield, and the old thermal linkage.
      • To access the cold head side of the linkage, we had to remove the lower conflat of the T.
    • We installed a 6" conflat flange, double faced with 4" bore (Lesker p/n DFF600X400), between the cryocooler and the T. This spacer raised the bottom face of the cold head to the correct height, so that the rigid copper bar runs approximately through the center of the shield apertures.
      • This required an order of new bolts with 3" length, to squeeze the three flanges (cryocooler, spacer, T) that are now stacked together.
    • We bolted the rigid copper bar to the coldhead, yawing the cryocooler to match the conflat bolt hole orientation while also pointing the copper bar down the axis of the arm.
      • This used new vented screws (UC Components p/n C-412-A) which are also silver plated, except in the location borrowed for the RTD stack as mentioned above [Attachment 7]. The RTD stack included a nut to adjust spring compression independently from screw threading.
      • Apiezon N thermal grease was applied on both surfaces to improve thermal conductivity across this joint.
      • We initially forgot to reinstall the mylar sheet radiation shielding that had been removed from the area around the cold head and around the linkage. This required that we reopen the bottom conflat to install the coldhead mitten, and that we pitch the aluminum shields away from the rigid bar to allow the mylar sheet to be inserted from the inside.
      • We found that the coldhead RTD had failed during intial mounting efforts, and a new RTD (actually one that had been desoldered from the setup previously, removed from the workpiece on 2021.08.07 but found to have no issue) was soldered and attached to the cold head, under the spring clamp.
    • Each shield was reinstalled, including:
      • The newly cut slots were used to pass shields over the rigid copper bar.
      • Electrical cabling was threaded through the usual apertures.
      • The outer shield was positioned on the G10 spacers.
      • Rough alignment was completed based on clearance from the rigid copper bar, and line of sight to optical window.
    • Final touchups were implemented:
      • Aluminum foil covered unused outer shield apertures.
      • A small aluminum foil panel was placed underneath the rigid copper bar, to cover the slot in the inner shield.
      • Final clamping of the inner shield and the heater (with indium gasket) were completed.
      • Thermal strap was used to link the cold baseplate to the rigid copper bar, with a bolted joint at the copper bar and a dog clamp joint at the baseplate. Apiezon N grease was applied to all contact faces.

The rest of the installation effort is captured in the next log post QIL/2695, to partition the items relevant to the radiative cooling of the silicon mass.

The photos here (and others) are posted to the QIL Cryo Vacuum Chamber photo album.

Attachment 1: IMG_0352.JPG
IMG_0352.JPG
Attachment 2: IMG_0353.JPG
IMG_0353.JPG
Attachment 3: IMG_0371.JPG
IMG_0371.JPG
Attachment 4: IMG_0370.JPG
IMG_0370.JPG
Attachment 5: IMG_0343.JPG
IMG_0343.JPG
Attachment 6: IMG_0351.JPG
IMG_0351.JPG
Attachment 7: IMG_0358.JPG
IMG_0358.JPG
Attachment 8: IMG_0346.JPG
IMG_0346.JPG
Attachment 9: IMG_2734.jpeg
IMG_2734.jpeg
Attachment 10: IMG_2740.jpeg
IMG_2740.jpeg
  2693   Fri Nov 12 10:54:55 2021 ranaSummary2um PhotodiodesResults from JPL PD: A1-Test3

IT would also be good if you could plot the RMS noise around 10 Hz and 100 Hz as a function of the bias and temperature, so we can see what the trends are. And how about post the data and scripts to the elog so we can munge data later?

  2692   Fri Nov 12 08:21:22 2021 AidanSummary2um PhotodiodesResults from JPL PD: A1-Test3

[Aidan]

Here is the analyzed data from Test 3 of the A1 JPL PD.

Note about the QE measurement: 

  • HOM beam QE was performed during the warm up phase. Collimating lens was fixed and beam pointing was optimized at 100mA before each measurement. Likely that not all power was on PD but that distribution was constant throughout measurement. Therefore, good proxy for shape of QE response.
  • Manual QE was performed with 25mA current, optimized collimating lens position (and thus the beam size on the PD). The data corresponds to 8.0mm between the lens mount and fiber mount. The beam pointing was optimized before each measurement at 25mA.
  • QE projection scales the "HOM beam QE" result to the manual QE measurements to project out expected QE performance vs temperature

Dark current is the output from the Keithley scan - the vertical scale is correct in Amps [ignore question mark]

Dark noise spectra are included for different bias levels and at different temperatures. Stll need to add ADC noise floor for these plots.

Notes from Test 3

~/JPL_PD/data/A1_test3/README

6-Oct-2021: done with cryocooler off and temperature increasing

PREAMP GAIN = 1E3

SR560 gain = 500

LD temp set point = 20.2kOhm

 

 

 

Attachment 1: A1_test3_nominal_QE.pdf
A1_test3_nominal_QE.pdf
Attachment 2: A1_test3_darkcurrent.pdf
A1_test3_darkcurrent.pdf
Attachment 3: A1_test3_noise_spectra_rev.pdf
A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf
  2691   Wed Oct 27 10:18:33 2021 Aidan, StephenSummaryCryo vacuum chamberInspection of Megastat post 408K event and pumping timeline

[Aidan]

I've attached a timeline of our inspection this afternoon along with today's pumping timeline,

Here is a brief summary of observations from previous pumping timelines. Today's pump down is consistent with previously observed timelines.


 

2:17PM – Assessing the impact of the 408K event in the MegaStat

Coldhead reached 363K (90C)

Innershield reached 403K (130C)

 

2:24PM – Aquadag E service temperature (149C)

https://www.laddresearch.com/index.php/lanotattachments/download/file/id/40/store/1/

Maximum service temperature in air* : 300°F (149°C) 

*Service temperature under vacuum conditions is significantly higher. Contact Acheson for specifics. 

2:25PM – Bringing Megastat back up to air for initial inspection

2:37PM – chamber is at air

 

2:41PM – removing bolts

 

 

3:13PM – initial inspection looks normal. No elevated amount of black particulates found on surfaces – consistent with or less than the amount seen last time we opened.

Stephen detected faint smell different from last time (“campfire”?)

  • One RTD connector did delaminate Aquadag from the inner shield

3:14PM - Stephen reattaching the RTD wiring that had delaminated. I wiped up visible particulates with isoproponal soaked wipe.

 

3:21PM – putting lid back on

3:30PM – lid on. Screws in finger tight

 

3:35PM – screws tight – ready to pump

3:40PM – pumping station on

 

Time (minutes) Pressure (Torr) Notes
0 7.5E2 Pirani gauge initially
1 7.5E2  
2 7.5E2  
4 7.5E2  
5 7.5E2  
7 7.5E2  
9 7.5E2  
10 7.5E2  
11 7.5E2  
12 4.3E2 Gauge starts reading decrease
13 2.2E2  
14 9.8E1  
15 5.5E1 Turbo ON
16 2.9E1  
17 1.6E1 Turbo at 44%
18 9.4E0  
19 4.9E0  
20 1.8E0 Turbo at 58%
21 3E-2 Turbo at 70%
22 1.1E-3 ION gauge readings from here. Turbo at 91%
23 7.5E-4 Turbo at 100%
24 6.3E-4  
25 5.9E-4  
26 5.5E-4

 (Cryo-cooler normally turned on around this time)

Not in this instance though

27 5.0E-4  
28 4.59E-4  
29 4.26E-4  
30 4.05E-4  
     

 

Attachment 1: IMG_5407.jpg
IMG_5407.jpg
Attachment 2: IMG_5408.jpg
IMG_5408.jpg
Attachment 3: IMG_5409.jpg
IMG_5409.jpg
Attachment 4: IMG_5410.jpg
IMG_5410.jpg
Attachment 5: IMG_5411.jpg
IMG_5411.jpg
Attachment 6: IMG_5412.jpg
IMG_5412.jpg
Attachment 7: IMG_5414.jpg
IMG_5414.jpg
Attachment 8: IMG_5415.jpg
IMG_5415.jpg
  2690   Tue Oct 26 08:43:38 2021 AidanUpdateCryo vacuum chamberCTC100 temperature alarm and heater shutoff

Instructions on how to enable the alarm and heater shut off for the CTC100.

Status: This reports the status of the alarm. If LATCH is enabled, this must be manually set to OFF once it has been enabled.

Mode: Set to "Level"

Latch: Optional to set to "YES" if desired.

Output: Set to "Heater"

Max: Set to desired maximum temperature.

The attached photos show:

  • the menu where the settings are ALARM entered
  • the main display just before the alarm is enabled (at 300.350K with a 1s delay)
  • the main display just after the alarm is enabled - note that the Heater Output has been set to 0W.
Attachment 1: Screenshot_from_2021-10-26_08-42-57.png
Screenshot_from_2021-10-26_08-42-57.png
Attachment 2: IMG_5371.jpg
IMG_5371.jpg
Attachment 3: IMG_5396.jpg
IMG_5396.jpg
Attachment 4: IMG_5401.jpg
IMG_5401.jpg
  2689   Tue Oct 26 07:32:52 2021 AidanUpdateCryo vacuum chamberHeater left on - chamber got warm

We're at 300K as of 7AM this morning.

Quote:

The cryocooler was switched off last Thursday to do testing on the JPL_PD. I turned the heater back on during this testing and neglected to turn it off when I finished at the end of the day. As a result, the workpiece reached ~400K over the weekend.

We are now allowing it to slowly cool down.

The CTC100 has a feature to specify an upper limit on temperature and then shut off the heater if that temperature is exceeded. We should engage this going forward.

 

  2687   Mon Oct 25 16:13:47 2021 AidanUpdateCryo vacuum chamberHeater left on - chamber got warm

The cryocooler was switched off last Thursday to do testing on the JPL_PD. I turned the heater back on during this testing and neglected to turn it off when I finished at the end of the day. As a result, the workpiece reached ~400K over the weekend.

We are now allowing it to slowly cool down.

The CTC100 has a feature to specify an upper limit on temperature and then shut off the heater if that temperature is exceeded. We should engage this going forward.

Attachment 1: Screenshot_from_2021-10-25_16-19-39.png
Screenshot_from_2021-10-25_16-19-39.png
  2686   Mon Oct 25 16:10:51 2021 AidanUpdateCDSRestarted computers and front-end model following campus power outage

Rebooted the workstations and FB4.

 Restarted the model on the FB4:

  • sudo /opt/rtcds/caltech/c4/target/c4iop/scripts/startupC4rt
  • sudo /opt/rtcds/caltech/c4/target/c4tst/scripts/startupC4rt
  2685   Fri Oct 22 13:31:36 2021 ChrisLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

Somewhere around the labs there should be a DataRay Beam'R2 scanning slit profiler, with an extended InGaAs detector that works out to 2.5 µm.

Quote:

agreed. You should put into Voyager chat and cryo/ET slack questins about 2 um beam profiling. We'll want it for anything 1.8-2.1 um.

Quote:

Definitely. I think the lack of beam profiling/imaging equipment is something we want to address too. We will waste a lot of time in Mariner if we can't profile our beams.

 

  2684   Fri Oct 22 12:04:13 2021 ranaLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

agreed. You should put into Voyager chat and cryo/ET slack questins about 2 um beam profiling. We'll want it for anything 1.8-2.1 um.

Quote:

Definitely. I think the lack of beam profiling/imaging equipment is something we want to address too. We will waste a lot of time in Mariner if we can't profile our beams.

  2683   Fri Oct 22 09:20:13 2021 AidanLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

Some of the data recorded during the current/micrometer scanning yesterday.

  • Distance between fiber/lens housings = Micrometer + 22.9mm
  • QE = Response / (1000Ohm*9.3E-4W/V) *(h*c/[e*L]) = Response * 0.667

Highlighted change to 25mA and also highest QE.

Time Temperature (K) Laser current (mA) Micrometer (mm) Peak Response (JPL/REF) QE
2:07PM 131.4 30 9.10 1.070 71.0%
2:14PM 133.5 30 9.10 1.090 73.0%
2:20PM 134.8 30 9.00 1.100 75.0%
2:25PM 136.4 30 8.80 1.134 75.6%
2:29PM 137.6 30 8.70 1.140 76.0%
2:31PM 138.1 30 8.60 1.120 74.7%
2:35PM 138.8 25 8.60 1.288 85.9%
2:42PM 140.5 25 8.80 1.246 83.1%
2:45PM   25 8.60 1.285 85.7%
2:249PM 142.2 25 8.40 1.310 87.4%
2:53PM 143 25 8.20 1.322 88.0%
3:01PM 144 25 8.00 1.328 88.6%
3:10PM   25 8.00 1.334 89.0%
3:14PM 147.5 25 7.80 1.314 87.6%
3:16PM 148.05 25 7.60 1.316 87.8%
3:19PM   25 8.00 1.328 88.6%
3:23PM 149.3 25 8.00 1.320 88.0%
3:30PM 151 25 8.00 1.315 87.7%
3:47PM 154 25 8.00 1.315 87.7%
4:02PM 157 25 8.00 1.300 86.7%
4:14PM 159.6 25 8.00 1.305 87.0%
  2682   Thu Oct 21 17:48:33 2021 AidanLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

Definitely. I think the lack of beam profiling/imaging equipment is something we want to address too. We will waste a lot of time in Mariner if we can't profile our beams.

Quote:

its worth looking into how fiber optic mode cleaning actually works:

https://doi.org/10.1201/9780203739662

In order to get a lot of cleaning you have to have a clean beam to begin with. There's a way to pre-clean by putting the laser output into a pinhole before coupling int othe single-mode fiber. Also, use a ~40-50m fiber to make sure the mod-mistanatched beam actually goes into the cladding rather than recombine into the Gaussian beam.

 

  2681   Thu Oct 21 16:52:26 2021 ranaLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

its worth looking into how fiber optic mode cleaning actually works:

https://doi.org/10.1201/9780203739662

In order to get a lot of cleaning you have to have a clean beam to begin with. There's a way to pre-clean by putting the laser output into a pinhole before coupling int othe single-mode fiber. Also, use a ~40-50m fiber to make sure the mod-mistanatched beam actually goes into the cladding rather than recombine into the Gaussian beam.

  2680   Thu Oct 21 15:32:32 2021 Aidan, RadhikaLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

[Aidan, Radhika]

We turned off the heater and the cryocooler this morning to around 11:30AM when the temperature of the diode was around 123K and are letting it gradually warm up.

Through the next 30K, we experimented with different distances between the fiber output and the collimating lens. Bias voltage always set to 1000mV. Laser diode current was set manually on the controller (the input from the DAC was unplugged as this is a little noisy). 

  1. We increased the distance between collimator and the lens by 1.4mm (from stage reading of 9.41mm to 10.8mm) and there was a small increase in the response (PD output/REF PD output)
  2. For each new setting, we set the laser diode current to 100mA and run the maximize script on the piezo mirror, adjusting the alignment to maximize the output power.
  3. We then stepped up the laser diode current in steps of 10mA from 25mA to 95mA and one last measurement at 101mA. The PD response dropped by 30-40% through this range.

It looked like the optimum translation stage setting was 9.75mm - however, i discovered something very interesting ...

If you run the maximize power script at 100mA to the laser diode, then drop the laser current to 30mA and rerun the script, you find that there is a different optimum alignment. This means that the output beam shape/pointing is power dependent. In other words, the output of the fiber is not properly mode-cleaned by the 2m patch cord we have.

Switching to 25mA, I optimized the alignment and continued exploring the optimum translation stage position. Dropping the stage position to 8.0mm maximized the response (at 25mA). Note that the code maximizes an EPICS channel called C4:TST-PD_RESPONSE which is (JPL_PD - DarkV)/REFPD. The reference PD filter bank has an offset applied so OUT16 has a mean value of zero when the laser is off. DarkV for the JPL PD is the PD voltage when the laser is off and this was manually updated every 5-10 minutes or so. C4:TST_PD_RESPONSE is not stored but the JPL_PD and REFPD channels are stored in frames.

On 27-Sept, I measured the ratio of power incident on the JPL PD to the voltage output from the REF PD: dP/dV = 9.3E-4 W/V. The JPL PD DC photocurrent sees a transimpedance gain of 1000Ohm. Therefore, QE is calculated using the following formula:

QE = (RESPONSE)*(1E-3/9.3E-4)*  h*c/(e*lambda) = RESPONSE*0.667

Using this calculation and a peak response value of 1.334 at about 145K, the peak QE was estimated to be about 89%. An error analysis is needed on this. And we need to figure out how to get a better output beam shape from the optical fiber (use a really long fiber?)

Note that the translation stage reading of 8.0mm corresponds to a fiber holder to lens mount distance of 30.9mm

 

Attachment 1: IMG_5305.jpg
IMG_5305.jpg
  2679   Mon Oct 18 15:25:14 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1

Initial running of analysis code puts the max QE at ~62 + /- 1% around 130-150K. I want to explore this temperature regime manually and see if we're saturating the PD or not.

3:30PM - Chamber is still under vacuum. Cryocooler turned back on.

Quote:

Terminated the data taking at 8:35Am this morning. The termperature traces of the cryo chamber show a couple of discontinuities in the gradient. I don't know what the cause is,

Quote:
  • We're at 164K as of 8AM this morning.

 

 

  2678   Mon Oct 11 08:35:21 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1

Terminated the data taking at 8:35Am this morning. The termperature traces of the cryo chamber show a couple of discontinuities in the gradient. I don't know what the cause is,

Quote:
  • We're at 164K as of 8AM this morning.

 

Attachment 1: Screenshot_from_2021-10-11_08-36-06.png
Screenshot_from_2021-10-11_08-36-06.png
  2677   Thu Oct 7 08:07:03 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1
  • We're at 164K as of 8AM this morning.
  2676   Wed Oct 6 13:50:18 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1
  • Turned cryocooler off around 1317588441 (about 1:46PM)
  • Restarted measurement with output going to JPL_PD/data/A1_test3
  • Room is noticeably quieter without the cryocooler on.
Quote:
  • Output going to JPL_PD/data/A1_test2 and DAQ
  • Test commenced at 8:20AM and cryo cooler started shortly afterwards
  • Once trhough the loop takes about 20 minutes
  • Cryocooler on at 8:42AM

 

  2675   Sun Oct 3 08:22:49 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1
  • Output going to JPL_PD/data/A1_test2 and DAQ
  • Test commenced at 8:20AM and cryo cooler started shortly afterwards
  • Once trhough the loop takes about 20 minutes
  • Cryocooler on at 8:42AM
  2674   Sat Oct 2 23:57:50 2021 StephenDailyProgressCryo vacuum chamberChamber pumping down, carbon paint flakes cleaned up

Pump down sequence executed tonight; Aidan plans to automate data collection during cooldown and warmup both, and the script will be activated early in the coming week.

- Carbon paint flakes (mentioned by Aidan in QIL/2667) were either picked up or scrubbed by IPA wipe, except the biggest, which were nudged near the closest 1/4-20 hole and picked up with tweezers for removal. There are still some small flakes of paint as there was less benefit to cleaning flakes closer to the PD or lens, and I opted not to risk any bumps. Aidan was correct, some areas of inner shield ID wall have flaked, but it seems the main location of delaminated paint is on the wrinkly foil excess covering the rim of the shield, an accidental paint location.

- While adding the lids, I monitored the PD outputs using Aidan's strip tool kindly left running. I never noticed any clipping in the trends - should I have been more skeptical?

- While adding the lids, I forgot to monitor the RTD outputs to the CTC100 controller, and the outer shield ended up shorting and ceasing to read any temperature. Didn't notice until I had turned on the roughing pump! Had to reopen and fix the short. Good reminder.

- Turbo came on at 12:07 am on 03 October at a pressure of 30 Torr - the setting is actually a timer and not a gauge reading.

  2673   Fri Oct 1 17:41:25 2021 PacoMiscEquipment transferFiber alignment pen

We borrowed the fiber alignment pen from the QIL for quickly coupling the AOM first-order beam in the ECDL experiment in DOPO lab.

  2672   Tue Sep 28 08:23:11 2021 AidanUpdate2um PhotodiodesDistances between optics, collimating lens, focussing lens and photodiode

Precise distances required between:

  • fiber launcher and collimating lens
  • focussing lens and beam waist 

accounting for thickness of optic mounts, sunken fiber launcher plane, back focal length of lenses, dispersive variation in focal lengths of lenses  from nominal and distance between PD surface and base of PD mount. Also shown are the distances between the steering mirrors (PZT steering mirror, lower periscope mirror and upper periscope mirror).

Beam propagation through this system is shown in the attached PDF. The upper plot shows a paraxial beam propagation as the collimating lens is displaced from the nominal position. The purpose is to indicate the beam size (radius) all the way through the system. We would like this to be less than about 6mm radius (12mm diameter) on all of our 1 diameter optics. The second plot shows the waist size at the PD as the collimating lens is moved by +/- 2mm. The purpose is to allow us to tune the beam size on the PD without clipping the beam on intervening optics. 

Keeping the collimating lens Delta Z to a range of +/- 2mm is safe for beam propagation in terms of clipping on apertures or on the 1.5mm diameter PD.

Attachment 1: JPL_PD_collimating_lens_Optical_layout.pdf
JPL_PD_collimating_lens_Optical_layout.pdf
Attachment 2: JPL_PD_optical_propagation_and_beam_size.pdf
JPL_PD_optical_propagation_and_beam_size.pdf
  2671   Wed Sep 22 16:40:19 2021 AidanLaser2um PhotodiodesBeam size measurements of the 2um beam on the PD

I performed some occlusion measurements of the 2um laser going into the cryo chamber. For different values of dz on the collimating lens translation stage, I moved the power meter into the beam using it's translation stage by an amount dx.

One the beam was on the power meter (aperture = 5mm diameter) the power stayed constant for several MM before dropping again (indicating all the laser beam was on the power meter).

There was a big inrcease in incident power as dz was increased. This, and the constant power across the PD aperture, indicates that the beam is clipping or sees an aperture somewhere like the focussing lens (f=75mm) or further upstrean. I will review the expected beam size as a function of position, assuming the given NA fof the fiber.

 

Attachment 1: PD_measurement_layout.png
PD_measurement_layout.png
Attachment 2: PD_occlusion_measurements.txt
	Length from Focussing lens to POW METER = 80mm															
																
																
2.73mm			3.73mm			dz	4.73mm	5.73mm	6.73mm	7.73mm	8.73mm	9.73mm	10.73mm	11.73mm	12.73mm	13.73mm
																
dx (0.01mm)	PM(uW)		dx	PM(uW)		dx	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)
60	191		40	3		200	241.9	272	305	348	395	454	523	605	697	798
50	189		50	7.25		175	241.7	270	307	346	394	452	519	595	685	777
35	181.7		65	23.6		150	234.5	262	296	333	377	432	492	563	645	727
20	169		75	85.5		130	218.8	243.6	274	308	356	395	445	507	579	657
... 7 more lines ...
  2670   Mon Sep 20 14:26:38 2021 ranaComputing2um PhotodiodesAutomation and analysis scripts for 2um data taking

you can put these in the GIT repo for the QIL Cryo tests that Radhika set up. Otherwise, they'll get lost. And we should probably change autorun to a .py script and document these in the README on the repo.

Quote:

The attached files are the scripts used to take data during the PD temperature cycling/testing and to retrieve and analyze data after the fact.

  • ~/JPL_PD/scripts/autorun2021.sh
    • ~/JPL_PD/scripts/piezo_mirror/maximize_output_power.py
  • ~/JPL_PD/data/A1_analysis/A1_analysis.py

 

  2669   Thu Sep 16 10:33:59 2021 AidanComputing2um PhotodiodesAutomation and analysis scripts for 2um data taking

The attached files are the scripts used to take data during the PD temperature cycling/testing and to retrieve and analyze data after the fact.

  • ~/JPL_PD/scripts/autorun2021.sh
    • ~/JPL_PD/scripts/piezo_mirror/maximize_output_power.py
  • ~/JPL_PD/data/A1_analysis/A1_analysis.py
Attachment 1: autorun2021.sh
#diode name
i=1001
diode=A1
caput C4:TST-FM15_OFFSET 0
sleep 1
while :; do
        #-----------------------------------------------------
        # dark current
        echo =======================
        echo ----- TOP OF LOOP -----
... 141 more lines ...
Attachment 2: maximize_output_power.py
# script to maximize the output power of the piezo
import serial
import time
import os, sys, subprocess
import numpy as np

def slowDownJog(ser):
    ser.write('1SU50\r\n')
    time.sleep(0.1)
... 195 more lines ...
Attachment 3: A1_analysis.py
# analysis od the A1 JPL PD diode
# Aidan Brooks - 10-Sept-2021

import cdsutils
import numpy as np
import matplotlib.pyplot as plt
import os, glob
import scipy.signal

... 172 more lines ...
  2668   Wed Sep 15 08:33:39 2021 AidanUpdate2um PhotodiodesVideo review of 2um testing setup post A1 testing

https://dcc.ligo.org/LIGO-G2102040

 

  2667   Wed Sep 15 08:22:32 2021 AidanDailyProgressCryo vacuum chamberCONTAMINATION: Black paint flecks throughout chamber

I was setting up for some characterization measurements of the JPL PD and I noticed that there are flecks of black paint all through the chamber. There were a couple of visible bare sections on the wall of the inner shield where paint had been removed.

Quote:

Monday I completed the vent that Aidan had started by turning off the cryocooler. During the afternoon I turned off the pumps, unbolted the chamber lid, and removed the radiation shield lids.

Next, Aidan was going to run some characterization measurements and determine whether to swap the diode or repeat with A1.

 

Attachment 1: IMG_4686.jpg
IMG_4686.jpg
Attachment 2: IMG_4687.jpg
IMG_4687.jpg
Attachment 3: IMG_4689.jpg
IMG_4689.jpg
Attachment 4: IMG_4690.jpg
IMG_4690.jpg
ELOG V3.1.3-